OpenStudy (anonymous):

Calculus: Evaluate lim as x->1 of (1- (1/x))/(1-(1/sqrt(x)))

OpenStudy (anonymous):

\[\large \frac{1-\frac{1}{x}}{1-\frac{1}{\sqrt{x}}}\]?

OpenStudy (anonymous):

Yes @satellite73

OpenStudy (anonymous):

Is the answer is 2?

OpenStudy (anonymous):

According to wolfram alpha, yes the answer is 2, but please help me how to approach this problem?

OpenStudy (anonymous):

idk i didn't to it the first obvious thing to do since it is of the form \(\frac{0}{0}\) is to use l'hopital, but that is probably no allowed and it is not necessary the next thing would be to get rid of the compound fraction

OpenStudy (anonymous):

or use wolfram lol

OpenStudy (anonymous):

Nope l'hopital at this point is no good haha

OpenStudy (anonymous):

can you think of anything to sandwich this with?

OpenStudy (anonymous):

k then get rid of the compound fraction

OpenStudy (anonymous):

no sandwich, it is all algebra

OpenStudy (anonymous):

First,simplify the numerator and denominator.

OpenStudy (anonymous):

okay.. I'll try getting rid of the compound fraction then

OpenStudy (anonymous):

multiply top and bottom by \(x\)

OpenStudy (anonymous):

then probably it would be a good idea to rationalize the denominator

OpenStudy (anonymous):

Rationalise the denominator.

OpenStudy (anonymous):

isnt the second step\[\frac{x-1}{x-\frac{x}{\sqrt(x)}}\] ?

OpenStudy (anonymous):

good question! yes but \[\frac{x}{\sqrt{x}}=\sqrt{x}\]

OpenStudy (anonymous):

oh wow.. I learn new things everyday haha, thanks for that

OpenStudy (anonymous):

you are four steps away 1) rationalize the denomiator 2) factor the denominator 3) cancel (which was the whole point of this) 4) plug in 1

OpenStudy (anonymous):

ah you knew \[\frac{x}{\sqrt{x}}=\sqrt{x}\] already either subtract the exponents, or else square \(\frac{x}{\sqrt{x}}\) and see that you get \(x\)

OpenStudy (anonymous):

did you complete step one yet? rationalize the denominator?

OpenStudy (anonymous):

I've actually successfully evaluated the limit. Thanks so much!!

OpenStudy (anonymous):

And no it wasn't obvious at first that x/sqrt(x) is just sqrt(x)

OpenStudy (anonymous):

yw hope it was clear gimmick is to cancel \(x-1\) to get rid of the 0 top and bottom

OpenStudy (anonymous):

I had to evaluate it to see for myself haha. thanks!